# Frank–Wolfe algorithm for DC optimization problem

### v2

@@ 3 Piecewise-star-convex functions
In the piecewise setting, $X_{i}^{*}$ consists of the global minimizers of $f$ restricted to the cell ${\cal C}_{i}$, which are in general only local minimizers on the whole set $\cal C$}. Corollary 3.12 refines the global statement for piecewise-star-convex functions by showing that Clarke stationarity implies cellwise optimality. Moreover, if the cell ${\cal C}_{i_{0}}$ containing $\bar{x}$ attains the global minimum value, i.e., $\min_{{\cal C}_{i_{0}}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}. This highlights the flexibility of piecewise-star-convexity: by allowing the domain to be partitioned into regions where the function behaves “locally convex,” one captures a broader class of DC objectives than with classical star–convexity, while still retaining descent inequalities sufficient to establish convergence of Frank–Wolfe variants.
4 The DC optimization problem
$\begin{array}[]{c}\min_{x\in{{\cal C}}}f(x):=g(x)-h(x),\end{array}$ (14)
  • (A)
    $\|\nabla g(x)-\nabla g(y)\|\leq L\|x-y\|$ for all   $x,y\in{{\cal C}}$.
Since ${{\cal C}}\subset{\mathbb{R}}^{n}$ is a compact, the study of problem (14) is bounded from below. Then, the optimum value of the problem (14) satisfies $\infty<f^{*}:=\inf_{x\in{{\cal C}}}f(x)$ and optimal set $\cal C$$^{*}$ is non-empty. According to Theorem 2.3, the first-order optimality condition for problem (14) is stated as
$(\nabla g({\bar{x}})-{\bar{u}})^{T}(x-{\bar{x}})\geq 0,\qquad\forall{\bar{u}}\in\partial g({\bar{x}}),\quad\forall x\in{{\cal C}}.$ (16)
In general, the condition (16) is necessary but not sufficient for optimality. A point ${\bar{x}}\in{{\cal C}}$ satisfying condition (16) is called a stationary point to problem (14). Consequently, all $x^{*}\in{{\cal C}}^{*}$ satisfies (16). We finish this section by presenting a variant of a classical example of DC optimization problem that fits the aforementioned requirements.
Example 4.1.
Let ${\cal C}_{i}\subset{\mathbb{R}}^{n}$ be an arbitrary set, for all $i=1,\ldots,m$. The square distance with respect to the set ${\cal C}_{i}$, denoted by $d^{2}_{{\cal C}_{i}}:{\mathbb{R}}^{n}\to{\mathbb{R}}$ is defined
$d^{2}_{{\cal C}_{i}}(x):=\inf_{y\in{\cal C}_{i}}\|x-y\|^{2}.$ (17)
In general, the function $d^{2}_{{\cal C}_{i}}$ is not convex. However, due to $\|x-y\|^{2}=\|x\|^{2}-2x^{T}y+\|y\|^{2}$, the distance function $d^{2}_{{\cal C}_{i}}$ becomes to
$d^{2}_{{\cal C}_{i}}(x):=\|x\|^{2}-\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right),$ (18)
which is a difference of two convex functions, see [23, Example 2.1.4]. Indeed, $d^{2}_{{\cal C}_{i}}=g-h$, where $g(x)=\|x\|^{2}$ is a convex quadratic function and $h(x)=\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)$ is a convex function given by the supremum of affine functions $x\mapsto\ell_{y}(x):=2x^{T}y-\|y\|^{2}$, for $y\in{\cal C}_{i}$. Let us state the constrained generalized Fermat-Weber location problem with square distances. For that, take $\omega_{i}\geq 0$, $i=1,\ldots,m$, such that $\sum_{i=1}^{m}\omega_{i}=1$ and note that
$\sum_{i=1}^{m}\omega_{i}d^{2}_{{\cal C}_{i}}(x)=\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right).$
Thus, for a given constrained set compact and convex ${\cal C}\subset{\mathbb{R}}^{n}$, the DC version of the constrained generalized Fermat-Weber location problem is stated as follows
$\begin{array}[]{c}\min_{x\in{{\cal C}}}\left(\|x\|^{2}-\sum_{i=1}^{m}\omega_{i}\sup_{y\in{\cal C}_{i}}\left(2x^{T}y-\|y\|^{2}\right)\right).\end{array}$ (19)
Finally, observe that the objective function in problem (19) satisfies all the assumptions of problem 14. Moreover, in certain cases the objective function of the generalized Fermat–Weber problem (19) is weak–star–convex. For example, if $m=n$ and ${\cal C}_{i}=\{e_{i},-e_{i}\}$, where $e_{i}$ denotes the $i$-th canonical vector of $\mathbb{R}^{n}$, then the objective becomes $f(x)=\|x\|^{2}-2\sum_{i=1}^{n}\omega_{i}|x_{i}|+1,$ which, by the same reasoning as in Example 3.5, can be shown to be weak–star–convex.
5 Classical Frank–Wolfe algorithm with an adaptive stepsize
In this section, we present the classic Frank–Wolfe algorithm with an adaptive stepsize to solve problem (14) and provide its convergence analysis. Although we assume condition (A) for $g$, the Lipschitz constant will not be used explicitly to determine the stepsize. Instead, the stepsize is computed adaptively using the scheme introduced in [3, 5], which estimates the Lipschitz constant on the fly. Our convergence analysis shows that, for piecewise-star-convexfunctions, possibly non–differentiable and non–convex, the method achieves an ${\cal O}(1/k)$ convergence rate for both the function values and the duality gap. It is worth noting that a variant of the classic Frank–Wolfe algorithm for solving problem (14) was proposed in [26], where the stepsize depends on a curvature/Lipschitz–type constant of the gradient of $f$. For non–convex objective functions, the convergence analysis in [28] shows that such a variant attains only an ${\cal O}(1/\sqrt{k})$ rate in terms of reaching a stationary point.
@@ 5 Classical Frank–Wolfe algorithm with an adaptive stepsize
Throughout this section, we assume access to a linear optimization oracle (LO oracle) that can minimize linear functions over the set $\cal C$}. The algorithm is stated below.
@@
Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}(x_{k})$ as follows
$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}(x_{k}):=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (20)
@@
If ${\omega}(x_{k})=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}(x_{k})|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}(x_{k})|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (21)
@@
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}(x_{k})|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (22)
@@
Let us describe the main features of Algorithm 1. First of all, note that due to $x^{0}\in{C}$, $p^{k}\in{C}$ and $\lambda_{k}\in(0,1]$ for all $k=0,1,\ldots$, and C{C} be convex set, it follows from inductive arguments and first equality in (23) that $(x^{k})_{k\in{\mathbb{N}}}\subset{C}$. Note that (20) implies that
${\omega}(x_{k})\leq(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\forall p\in{C}.$ (24)
Consequently, if ${\omega}(x_{k})=0$, then $(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k})\geq 0$, for all $p\in{C}$, i.e., $x^{k}$ is a stationary point to problem (14). Note that letting $p=x^{k}$ in (24) we conclude that
${\omega}(x_{k})\leq 0,\qquad k=0,1,\ldots.$ (25)
In order to simplify the notations, from now on we will use the following two notations:
${\omega}_{k}:={\omega}(x_{k}).$ (26)
Since $\omega_{k}=0$ implies that $x^{k}$ is a stationary point to problem (14). Thus, in view of (25) from now on we assume that ${\omega}_{k}<0$, for all $k\in{\mathbb{N}}$. The next result shows that the backtracking line–search in Algorithm 1 is well defined, identifies the smallest $j_{k}$ in Step 4, and establishes a descent inequality for the sequence $(x^{k})_{k\in{\mathbb{N}}}$.
@@
In addition, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. And the following inequality holds
@@
Hence, it follows from the last inequality that (22) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proofs the first statement. Consequently, $j_{k}$ is well defined and is the smallest non-negative integer satisfying (27) and (28). Therefore, the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is well defined. We proceed with the proof of (29), by showing the following inequality
@@
For that we separately consider two cases: $\lambda_{k}=|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$ and ${j_{k}}=1$. In the former, we have $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}=|{\omega_{k}}|/{\lambda_{k}}$,which substituting into the left hand side of (32) yields an equality. If now $\lambda_{k}=1$, then it follows from (21) that $\lambda_{k}=1\leq|{\omega_{k}}|/({2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2})$ or equivalently ${2^{j_{k}}L_{k}}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|$ and $\lambda_{k}=1$, which again substituting into the left hand side of (32) yields the inequality. Hence, (32) holds. Therefore, (29) follows by combining (28) with (32), which concludes the proof. ∎
@@
The sequence $(L_{k})_{k\in{\mathbb{N}}^{*}}$ satisfies the following inequalities
$L_{0}\leq L_{k}\leq L+L_{0},\quad\qquad k=0,1,\ldots.$ (34)
and the step size sequence $(\lambda_{k})_{k\in{\mathbb{N}}}$ satisfies
$\lambda_{k}\geq\min\left\{1,\frac{|{\omega}_{k}|}{\alpha}\right\},\quad\qquad k=0,1,\ldots.$ (35)
@@
We proceed by induction to prove (34). Since the Lipschitz constant $L>0$, (34) holds for $k=0$. Now, assume that (34) holds for some $k>0$. We separately consider two cases: $j_{k}=0$ and $j_{k}\geq 1$. First assume that $j_{k}=0$. In this case, Step 1 together with second equality in (23) imply that $L_{k+1}=2^{-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}=0$. Since we are under the assumption that (34) holds, we have $L_{k+1}=2^{-1}L_{k}<L_{k}\leq L+L_{0}$. Thus, we conclude that two inequalities in (34) holds for $k+1$ and $j_{k}=0$. Next, we assume that $j_{k}\geq 1$. Again, Step 1 together with second equality in (23) yield $L_{k+1}:=2^{j_{k}-1}L_{k}\geq L_{0}$. Thus, the fist inequality in (34) holds for $k+1$ and $j_{k}\geq 1$. To prove the second one, we assume by contradiction that $L_{k+1}>L+L_{0}$. This condition and (23) imply that $2^{j_{k}-1}L_{k}>L$. Thus, the first part of Proposition 5.1 implies that $j=j_{k}-1$ satisfies (22). Since the condition $L_{k+1}>L+L_{0}$ together (23) also implies $2^{j_{k}-1}L_{k}>L_{0}$, we have a contradiction because Proposition 5.1 also implies that $j_{k}$ is the smallest non-negative integer satisfying (27) and (28). Therefore, (34) holds for $k+1$, which concludes the proof that (34) holds for all $k\in{\mathbb{N}}$. The inequality (35) is an immediate consequence of (15), (21), the second inequality in (34) and (33). ∎
Next, we state iteration–complexity bounds for the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1. Since the analysis closely follows that in [4], the proofs are omitted.
@@
$\lim_{k\to+\infty}{\omega}_{k}=0$. Consequently, every limit point of the sequence $(x^{k})_{k\in{\mathbb{N}}}$ generated by Algorithm 1 is a stationary point to problem (14). In addition, for every $N\in{\mathbb{N}}$, there holds
$\min\left\{|{\omega}_{k}|:~k=0,1,\ldots,N\right\}\leq\max\left\{{{2(f(x^{0})-f^{*})}},\sqrt{2\alpha{(f(x^{0})-f^{*})}}\right\}\frac{1}{\sqrt{N+1}}.$
Since $\cal C$}} is compact and the functions $g$ and $h$ are convex, we can define the following constants:
$\Theta:=\min\left\{\frac{2}{{\bar{\Gamma}}\operatorname{diam}({\cal C})},\frac{1}{\alpha}\right\}\qquad{\bar{\Gamma}}:=\max_{x\in{{\cal C}}}\{\|\nabla g(x)\|+\|u\|:~u\in\partial h(x)\},\qquad\beta:=\frac{{\bar{\Gamma}}^{2}}{{\Theta L_{0}}}.$
Since the proof of the next theorem follows a standard argument in the convergence analysis of line–search methods, we omit it.

### v3

@@ 4 Piecewise star-convexity with a nonsmooth DC structure
In the piecewise setting, $X_{i}^{*}=\arg\min_{{\cal C}_{i}}f$ consists of minimizers of $f$ restricted to the cell ${\cal C}_{i}$. Such points are global minimizers on ${\cal C}_{i}$, but they need not be minimizers of $f$ on the whole set $\cal C$} and may fail to be even local minimizers in $\cal C$}. This is precisely the situation in which piecewise star-convexity is useful, star-type inequalities are enforced within each cell with respect to $X_{i}^{*}$, even though they may break down across cell boundaries. Corollary 4.10 refines the global picture by showing that Clarke stationarity implies cellwise optimality, namely $\bar{x}\in X_{i}^{*}$ for the cell ${\cal C}_{i}$ containing $\bar{x}$. In particular, if that cell achieves the smallest cell minimum value, that is, if $\min_{{\cal C}_{i}}f=\min_{{\cal C}}f$, then $\bar{x}$ is a global minimizer of $f$ on $\cal C$}.
5 Frank–Wolfe algorithm
In this section, we present the Frank–Wolfe algorithm for solving (1) under a Lipschitz-based adaptive stepsize rule. We also establish iteration-complexity guarantees under piecewise star-convexity, see [3, 5]. Our analysis shows that, for piecewise star-convex functions with a nonsmooth DC structure, the method attains an ${\cal O}(1/k)$ rate for the function values (relative to a best cell minimum) and a corresponding ${\cal O}(1/k)$ bound for the Frank–Wolfe gap. For comparison, a related Frank–Wolfe variant for (1) was proposed in [23], with stepsizes based on a curvature/Lipschitz-type constant. For general nonconvex objectives, the analysis in [26] yields an ${\cal O}(1/\sqrt{k})$ rate in terms of stationarity.
@@ 5 Frank–Wolfe algorithm
Throughout this section, we assume access to a linear optimization oracle (LO oracle) that minimizes linear functions over $\cal C$}. The algorithm is stated below.
@@
Use an “LO oracle” to compute an optimal solution $p^{k}$ and the optimal value ${\omega}_{k}$ as follows
$p^{k}\in{\rm argmin}_{p\in{{\cal C}}}(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p-x^{k}),\qquad\quad{\omega}_{k}:=(\nabla g(x^{k})-u^{k})^{{\scriptscriptstyle\mathrm{T}}}(p^{k}-x^{k}).$ (16)
@@
If ${\omega}_{k}=0$, then stop. Otherwise, compute the step size $\lambda_{j}\in(0,1]$ as follows
$\lambda_{j}=\mbox{min}\left\{1,\frac{|{\omega}_{k}|}{{2^{j}L_{k}}\|p^{k}-x^{k}\|^{2}}\right\}:={\rm argmin}_{\lambda\in(0,1]}\left\{-|{\omega}_{k}|\lambda+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda^{2}\right\}.$ (17)
@@
$f(x^{k}+\lambda_{j}(p^{k}-x^{k}))\leq f(x^{k})-|{\omega}_{k}|\lambda_{j}+\frac{{2^{j}L_{k}}}{2}\|p^{k}-x^{k}\|^{2}\lambda_{j}^{2},$ (18)
@@
The LO oracle provides $p^{k}\in{\cal C}$ and induces the direction $d^{k}:=p^{k}-x^{k}$, with $p^{k}:=p(x^{k},u^{k})$ as defined in in (5). The scalar quantity $\omega_{k}:=\omega(x^{k},u^{k})\leq 0$, where $\omega(x^{k},u^{k})$ as defined in (5) is the corresponding Frank–Wolfe gap and vanishes only at stationary points, see Proposition 3.1. Therefore, the algorithm stops whenever $\omega_{k}=0$. From now on, we assume
$\omega_{k}<0,\qquad\qquad k=0,1,\ldots.$ (20)
for all $k\geq 0$, so the method generates an infinite sequence $\{x^{k}\}_{k\in\mathbb{N}}\subset{\cal C}$. In fact, since the update of $x_{k+1}$ in  (19) is a convex combination of $x^{k}$ and $p^{k}$ with $\lambda_{k}\in(0,1]$, the convexity of $\cal C$} yields $x^{k}\in{\cal C}$ for every $k$.
Next we apply Proposition 2.4 to show that the backtracking procedure in Step 2 and 3 terminates in finitely many trials at each iteration. Therefore, $j_{k}$, $\lambda_{k}$, and the update $L_{k+1}$ in (19) are well defined. The next proposition states this formally and records the key descent inequality used in the complexity analysis.
@@
In addition, the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 is well defined. And the following inequality holds
@@
Hence, it follows from the last inequality that (18) holds, for all $j\in{\mathbb{N}}$ such that ${2^{j}L_{k}}\geq L$, which proves the first statement. Consequently, $j_{k}$ is well defined and is the smallest non-negative integer satisfying (21) and (22). Therefore, the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 is well defined. We proceed with the proof of (23), by showing the following inequality
@@
For that, we note that $0\leq\lambda_{k}\geq|{\omega_{k}}|/{(2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2})}$, from which we obtain that $2^{j_{k}}L_{k}\|p^{k}-x^{k}\|^{2}\leq|{\omega_{k}}|/{\lambda_{k}}$. Substituting this into the left hand side of (26) yields the inequality. Hence, (26) holds. Therefore, (23) follows by combining (22) with (26), which concludes the proof. ∎
5.1 Convergence analysis
In this section we analyze Algorithm 1 under Assumption (A). We first show that the backtracking rule yields uniformly bounded Lipschitz estimates and a uniform lower bound on the accepted stepsizes. These bounds, combined with the sufficient decrease inequality in Proposition 5.1, imply monotone descent of $\{f(x^{k})\}_{k\in\mathbb{N}}$, $\lim_{k\to\infty}\omega_{k}=0$, and Clarke-stationarity of every accumulation point. If $f$ is piecewise-star-convex, any such limit point is a global minimizer on its cell and a local minimizer of $f$ whenever it lies in the interior of that cell.
@@ 5.1 Convergence analysis
The next lemma provides a uniform upper bound on the backtracking estimates produced, and yields a convenient lower bound on the accepted stepsizes $\lambda_{k}$.
@@
Assume (A) and let $(L_{k})_{k\in\mathbb{N}}$ be generated by (19). Then
$0<L_{k}\leq L+L_{0},\qquad k=0,1,\ldots.$ (28)
Moreover, the stepsizes $(\lambda_{k})_{k\in\mathbb{N}}$ selected by Lipschitz-based adaptive stepsize satisfy
$\lambda_{k}\ \geq\ \min\Bigl\{1,\ \frac{|{\omega}^{k}|}{\alpha}\Bigr\},\qquad k=0,1,\ldots.$ (29)
@@
We first prove (28). Since $L_{0}>0$, we have $0<L_{0}\leq L+L_{0}$. Assume that $0<L_{k}\leq L+L_{0}$ for some $k\geq 0$ and recall from (19) that $L_{k+1}=2^{\,j_{k}-1}L_{k}.$ If $j_{k}=0$, then $L_{k+1}=\tfrac{1}{2}L_{k}$, hence $0<L_{k+1}\leq L_{k}\leq L+L_{0}$. If $j_{k}\geq 1$, suppose by contradiction that $L_{k+1}>L+L_{0}$. Then $2^{\,j_{k}-1}L_{k}=L_{k+1}>L$, which implies that $2^{\,j_{k}-1}L_{k}\geq L$. By the first statement of Proposition 5.1, the backtracking inequality (18) holds for $j=j_{k}-1$. This contradicts the fact that $j_{k}$ is the first index accepted by the backtracking loop in Step 3.2 of Lipschitz-based adaptive stepsize procedure. Hence $L_{k+1}\leq L+L_{0}$, and in all cases $0<L_{k+1}\leq L+L_{0}$. By induction, (28) holds for all $k$. We now prove (29). By (17) with $j=j_{k}$,
$\lambda_{k}=\min\Bigl\{1,\ \frac{|\omega_{k}|}{2^{\,j_{k}}L_{k}\,\|p^{k}-x^{k}\|^{2}}\Bigr\}.$
Since $x^{k},p^{k}\in{\cal C}$, we have $\|p^{k}-x^{k}\|\leq\operatorname{diam}({\cal C})$ by (2). Moreover, from (19) we obtain $2^{\,j_{k}}L_{k}=2L_{k+1}$, and then (28) yields $2^{\,j_{k}}L_{k}=2L_{k+1}\leq 2(L+L_{0}).$ Combining these estimates gives
$\lambda_{k}\geq\min\Bigl\{1,\ \frac{|\omega_{k}|}{2(L+L_{0})\,\operatorname{diam}({\cal C})^{2}}\Bigr\}=\min\Bigl\{1,\ \frac{|\omega_{k}|}{\alpha}\Bigr\},$
where $\alpha$ is defined in (27). This is (29) and the proof is concluded. ∎
In the next theorem, we summarize the resulting convergence guarantees for Algorithm 1.
@@
Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let $\{x^{k}\}_{k\in\mathbb{N}}$ be generated by Algorithm 1. Then, the following statement holds:
  • (i)
    The objective values sequence $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing. Moreover, the Frank–Wolfe gap satisfies $\lim_{k\to\infty}\omega_{k}=0.$
  • (ii)
    Every accumulation point $\bar{x}$ of $\{x^{k}\}_{k\in\mathbb{N}}$ is Clarke-stationary for (1);
  • (iii)
    If, in addition, $f$ is piecewise-star-convex on $\cal C$} with respect to a partition ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$, then for any such limit point $\bar{x}$ and any $i_{*}\in I$ with $\bar{x}\in{\cal C}_{i_{*}}$ and $\bar{x}\in X_{i_{*}}^{*}$, i.e., $\bar{x}$ is a global minimizer of $f$ on its cell ${\cal C}_{i_{*}}$. Furthermore, if $\bar{x}\in{\rm int}\,{\cal C}_{i_{*}}$, where ${\rm int}\,{\cal C}_{i_{*}}$ denotes the interior of the set ${\cal C}_{i_{*}}$, then $\bar{x}$ is local minimizer of $f$.
Proof.
To prove item (i), first recall that $\omega_{k}\neq 0$ for all $k$, which implies that an infinite sequence is generated. By Proposition 5.1, the accepted stepsize satisfies the sufficient decrease inequality (23), and hence $0\leq\frac{1}{2}|{\omega}_{k}|\leq f(x^{k})-f(x^{k+1})$ for all $k=0,1,\ldots,$ which implies that $\{f(x^{k})\}$ is monotone nonincreasing. We proceed to prove that $\lim_{k\to\infty}\omega_{k}=0$. Let $\alpha>0$ be defined in (27). Combining the sufficient decrease (23) with the lower bound (29), we obtain
$\tfrac{1}{2}\min\bigl\{|{\omega}_{k}|,\tfrac{1}{\alpha}|{\omega}_{k}|^{2}\bigr\}\leq f(x^{k})-f(x^{k+1}),\qquad k=0,1,\ldots.$ (30)
Since $\{f(x^{k})\}_{k\in\mathbb{N}}$ is monotone nonincreasing and bounded below on $\cal C$}, it converges. Hence, we have $\lim_{k\to\infty}(f(x^{k})-f(x^{k+1}))=0$. Therefore, passing to the limit in inequality  (30) yields $\lim_{k\to+\infty}\min\bigl\{|{\omega}_{k}|,\tfrac{1}{\alpha}|{\omega}_{k}|^{2}\bigr\}=0,$ which implies that $\lim_{k\to\infty}\omega_{k}=0$.
To prove item (ii), let $\bar{x}$ be a limit point of $\{x^{k}\}_{k\in\mathbb{N}}$ and take a subsequence $\{x^{k_{\ell}}\}_{\ell\in\mathbb{N}}$ such that $\lim_{\ell\to+\infty}x^{k_{\ell}}=\bar{x}$. Because $h$ is convex and $\cal C$} is compact, Proposition 2.7 ensures that $\bigcup_{x\in{\cal C}}\partial h(x)$ is bounded. Hence, $\{u^{k_{\ell}}\}_{\ell\in\mathbb{N}}$ is bounded and, passing to a further subsequence if needed, we may assume $\lim_{\ell\to+\infty}u^{k_{\ell}}=\bar{u}$, for some $\bar{u}\in\mathbb{R}^{n}$. By Proposition 2.6, the graph of subdifferential $\partial h$ is closed, which implies that $\bar{u}\in\partial h(\bar{x})$. Now fix any $x\in{\cal C}$. Since $p^{k_{\ell}}$ minimizes the linear functional in (16), we have
$(\nabla g(x^{k_{\ell}})-u^{k_{\ell}})^{\top}(x-x^{k_{\ell}})\;\geq\;(\nabla g(x^{k_{\ell}})-u^{k_{\ell}})^{\top}(p^{k_{\ell}}-x^{k_{\ell}})=\omega_{k_{\ell}}.$
Using that $\lim_{\ell\to+\infty}x^{k_{\ell}}=\bar{x}$, $\lim_{\ell\to+\infty}u^{k_{\ell}}=\bar{u}$, the continuity of $\nabla g$, and $\lim_{\ell\to+\infty}\omega_{k_{\ell}}=0$, we obtain that $(\nabla g(\bar{x})-\bar{u})^{\top}(x-\bar{x})\geq 0$, for all $x\in{\cal C}$, which is precisely the Clarke-stationarity condition for (1).
Finally, we prove item (iii). Assume additionally that $f$ is piecewise-star-convex on $\cal C$} with respect to ${\cal P}=\{{\cal C}_{i}\}_{i\in I}$. Let $\bar{x}$ be any accumulation point of $\{x^{k}\}_{k\in\mathbb{N}}$. Choose an index $i_{*}\in I$ such that $\bar{x}\in{\cal C}_{i_{*}}$, such an index exists because $\cal P$} covers $\cal C$}. Since item (ii) has been proved, $\bar{x}$ is Clarke-stationary for (1). Therefore, Corollary 4.10 applies and yields $\bar{x}\in X_{i_{*}}^{*}$, i.e., $\bar{x}$ is a global minimizer of $f$ over ${\cal C}_{i_{*}}$. It remains to show the local minimality statement when $\bar{x}\in{\rm int}\,{\cal C}_{i_{*}}$. Because ${\rm int}\,{\cal C}_{i_{*}}$ is open in $\mathbb{R}^{n}$, there exists $r>0$ such that $B(\bar{x},r)\subset{\cal C}_{i_{*}}$. For every $y\in{\cal C}\cap B(\bar{x},r)$ we have $y\in{\cal C}_{i_{*}}$, and since $\bar{x}$ minimizes $f$ on ${\cal C}_{i_{*}}$ it follows that $f(\bar{x})\leq f(y)$, for all $y\in{\cal C}\cap B(\bar{x},r)$. Hence $\bar{x}$ is a local minimizer of $f$ on $\cal C$}. ∎
We end this section by emphasizing that, compared with the schemes in [4, 23], the only algorithmic ingredient that differs is the stepsize selection performed by the backtracking procedure in Algorithm 1. In the next section, we use the resulting descent estimate to derive the corresponding convergence and iteration–complexity guarantees.
5.2 Iteration-complexity analysis
In this section, we derive iteration–complexity bounds for the sequence $\{x^{k}\}_{k\in\mathbb{N}}$ generated by Algorithm 1 when applied to the DC problem (1). We focus on objectives of the form f=ghf=g-h that are piecewise star-convex on $\cal C$}, where $g$ is convex and continuously differentiable with $L$-Lipschitz continuous gradient on $\cal C$}, and $h$ is convex and possibly nonsmooth.
We begin with a structural lemma for finite convex partitions that requires only continuity of $f$ and monotonicity of the objective values. It states that, if the function values converge to the minimum attained on some cell of the partition, then after a finite number of iterations the sequence can only visit cells whose minimum value does not exceed that limit. This prefix–exclusion mechanism will be used later, once the descent analysis ensures that $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing.
Lemma 5.4.
Let ${\cal C}\subset\mathbb{R}^{n}$ be compact and convex, and let ${\cal P}=\{{\cal C}_{i}:i=1,\dots,m\}$ be a finite partition of $\cal C$} into nonempty convex sets. Let $f:\mathbb{R}^{n}\to\mathbb{R}$ be such that $f_{i}^{*}:=\min_{x\in{\cal C}_{i}}f(x)$ is well defined for every $i$. Let $\{x^{k}\}_{k\in\mathbb{N}}\subset{\cal C}$ be a sequence satisfying $f(x^{k+1})\leq f(x^{k})$ for all $k$, and suppose that, for some $i_{*}\in\{1,\dots,m\}$,
$\lim_{k\to\infty}f(x^{k})\;=\;f_{i_{*}}^{*}.$ (31)
Then there exists $N\in\mathbb{N}$ such that, for all $k\geq N$ and every index $i_{k}\in\{1,\dots,m\}$ with $x^{k}\in{\cal C}_{i_{k}}$,
$f_{i_{k}}^{*}\;\leq\;f_{i_{*}}^{*}.$
Proof.
Define $\hat{I}:=\{\,i\in\{1,\dots,m\}:\ f_{i}^{*}>f_{i_{*}}^{*}\,\}.$ If $\hat{I}$$\emptyset$, then the conclusion holds by taking $N:=0$. Assume $\hat{I}\neq\emptyset$ and fix $i\in\hat{I}$. By (31), since $f_{i}^{*}>f_{i_{*}}^{*}$, there exists $N_{i}\in\mathbb{N}$ such that $f(x^{N_{i}})<f_{i}^{*}.$ Because the sequence $\{f(x^{k})\}_{k\in\mathbb{N}}$ is nonincreasing, it follows that
$f(x^{k})\leq f(x^{N_{i}})<f_{i}^{*},\qquad\forall\,k\geq N_{i}.$
If, for some $k\geq N_{i}$, we had $x^{k}\in{\cal C}_{i}$, then by definition of $f_{i}^{*}$ as the minimum of $f$ on ${\cal C}_{i}$ we would have $f(x^{k})\geq f_{i}^{*}$, contradicting the strict inequality above. Hence,
$x^{k}\notin{\cal C}_{i},\qquad\forall\,k\geq N_{i}.$
Now set $N:=\max_{i\in\hat{I}}N_{i}$, finite because $\hat{I}$ is finite. Then for every $k\geq N$ we have $x^{k}\notin{\cal C}_{i}$ for all $i\in\hat{I}$. Therefore, if $x^{k}\in{\cal C}_{i_{k}}$ with $k\geq N$, necessarily $i_{k}\notin\hat{I}$, which means $f_{i_{k}}^{*}\leq f_{i_{*}}^{*}$, and the desired inequality is proved. ∎
We now state a complexity result showing that, despite nonconvexity, piecewise–star–convex objectives achieve the same Frank–Wolfe iteration bound as in the convex setting.